# Logarithmic or algebraic: roughening of an active Kardar-Parisi-Zhang surface

### v3

@@ Abstract
The Kardar-Parisi-Zhang (KPZ) equation sets the universality class for growing and roughening of nonequilibrium surfaces without any conservation law and nonlocal effects. We argue here that the KPZ equation can be generalised by including a symmetry-permitted nonlocal nonlinear term of active origin that is of the same order as the one included in the KPZ equation. Including this term, the 2D active KPZ equation is stable in some parameter regimes, in which the interface conformation fluctuations exhibit sub-logarithmic or super-logarithmic roughness, with nonuniversal exponents, giving positional generalised quasi-long-ranged order. For other parameter choices, the model is unstable, suggesting a perturbatively inaccessible algebraically rough interface or positional short-ranged order. Our model should serve as a paradigmatic nonlocal growth equation.
@@
The Kardar-Parisi-Zhang (KPZ) equation Kardar et al. (1986); Krug (1997); Barabási and Stanley (1995) for growing nonequilibrium surfaces displays a nonequilibrium roughening transition between a smooth phase, whose long wavelength scaling properties are identical to an Edward-Wilkinson (EW) surface Edwards and Wilkinson (1982) at the same dimension $d$, to a perturbatively inaccessible rough surface Barabási and Stanley (1995); Tang et al. (1990) when $d>d_{c}=2$, its lower critical dimension. Importantly, its local surface growth velocity that depends locally on surface fluctuations, and hence cannot describe nonequilibrium surface dynamics with nonlocal interactions.
@@
Theoretical studies on nonlocal interactions has a longstanding history in equilibrium systems Fisher et al. (1972); Frey and Schwabl (1994); Bayong et al. (1999); Aronovitz and Lubensky (1988); Banerjee et al. (2019a, b). Examples of their prominent nonequilibrium counterparts include interface dynamics involving nonlocal interactions, e.g., flame front propagation, thin film growth Pelcé et al. (2004). Monte-Carlo simulation studies on a solid-on-solid model with nonlocal effects Ketterl and Hagston (1999) reveal how the shading phenomena in surface growth is described by nonlocal effects. Furthermore, kinetic roughening in the presence of nonlocal interactions Mukherji and Bhattacharjee (1997) display generic non-KPZ scaling behaviour. Nonlocal effects are often important in biological growth processes; see, e.g., Ref. Santalla and Ferreira (2018) for a recent study. Furthermore in many applications, the growth is controlled by fast nonlocal transport, not included in the KPZ equation. Prominent examples include diffusion-controlled nonlocal transport Aharonov and Rothman (1996), dissolution or precipitation processes Lichtner et al. (1996), gas-solid reactions Doǧu (1981), and a variety of reaction engineering processes Delmono and Froment (1999). Yet another potentially relevant example is diffusion-limited erosion, that displays nonlocal stabilisation of surfaces Krug and Meakin (1991); see also Ref. Nicoli et al. (2009). Nonlocal effects may be vital even in geological contexts, e.g., earth surface roughness R. Schumer (2017). Inspired by these past studies, we explore the generic consequences of competition between local contributions and those that depend on the global surface profile, i.e., nonlocal contributions to the local surface velocity, by constructing a purpose-built conceptual model.
@@
In this Letter, we set up and study a generalisation of the KPZ equation, where the surface velocity depends, in contrast to the KPZ equation, nonlocally on the surface fluctuations. We do this by adding a symmetry-permitted nonlocal nonlinear gradient term that is of the same order as the usual KPZ nonlinear term. This allows us to study competition between local and nonlocal nonlinear effects. To generalise the scope of our study, we also include chiral contributions, which is ubiquitous in soft matter and biologically inspired systems; see, e.g., Refs. Martin et al. (2021); Lalitha et al. (2001). The resulting equation in 2D, named active-KPZ or a-KPZ equation is
@@
a nonlocal generalisation to the usual KPZ equation that is distinct from the one considered in Ref. Mukherji and Bhattacharjee (1997). Here, the tensor $e_{jm}$ is the 2D totally antisymmetric matrix. Further, $Q_{ij}({\bf r})$ is the longitudinal projection operator that in the Fourier space is $Q_{ij}({\bf k})=k_{i}k_{j}/k^{2}$, where ${\bf k}$ is a Fourier wavevector, and is nonlocal. Physically $\lambda_{1}Q_{ij}({\bf r})(\nabla_{i}h\nabla_{j}h)+\lambda_{2}Q_{ij}({\bf r})e_{jm}(\nabla_{i}h\nabla_{m}h)$ is the contribution to the surface velocity normal to the base plane $v_{p}=\partial h/\partial t$ that is nonlocal in height fluctuations ${\bm{\nabla}}h$. Noise $\eta$ is a zero-mean, Gaussian-distributed white noise with a variance $\langle\eta({\bf x},t)\eta(0,0)\rangle=2D\delta^{d}({\bf x})\delta(t)$. We extract the scaling of the stable phases, which exists for a range of the model parameters. In particular, we show that the variance $\Delta\equiv\langle h({\bf x},t)^{2}\rangle\sim[\ln(L/a)]^{\mu}$ for a surface of lateral size $L$, where $\mu<(>)1$ for sub-(super-)logarithmic roughness and $a$ is a microscopic cutoff. This defines positional generalised quasi-long-ranged order (QLRO), generalising the well-known QLRO of EW surfaces Barabási and Stanley (1995), in which $\Delta\sim\ln(L/a)$, i.e., $\mu=1$. Further, the time-scale of relaxation $\tau(L)\sim L^{2}\,(\ln\,L)^{-\kappa},\,\kappa>0$, i.e., logarithmically superdiffusive. Both $\mu$ and $\kappa$ are non-universal: They vary continuously with $\lambda_{1}/\lambda,\,\lambda_{2}/\lambda$.
The form of Eq. (15) can be obtained by first considering the mapping from the KPZ equation to the Burgers equation Forster et al. (1977) in terms of the “Burgers velocity ${\bf v}={\bm{\nabla}}h$”. Now generalise the Burgers equation nonlinearity $\lambda{\bm{\nabla}}v^{2}$ to $\lambda_{1}\nabla_{j}(v_{i}v_{j})+\lambda_{2}e_{jm}\nabla_{j}(v_{i}v_{m})$, and then writing them in terms of $h$ produces the $\lambda_{1}$- and $\lambda_{2}$-terms in (15); see Supplemental Material (SM).
The $\lambda_{1}$- and $\lambda_{2}$-terms in (15) can be motivated by considering a nearly flat nonequilibrium surface described by a single valued height field $h({\bf x},t)$ in Monge gauge  Chaikin and Lubensky (2000); Nelson et al. (1989), with an active conserved density $\rho({\bf x},t)$ living on it. Its hydrodynamic equation, retaining only the lowest order in nonlinearities and spatial gradients, reads
@@
where $v(\rho)$ is a local density-dependent velocity of the membrane; $v(\rho)=v_{0}+g_{1}\rho$ to the leading order in $\rho$; $g_{1}$ is a coupling constant of either sign. Further, density $\rho$ follows $\partial_{t}\rho=-{\bm{\nabla}}\cdot{\bf J}$, where $\bf J$ is the current. The specific form of the particle dynamics decides the structure of $\bf J$. We choose ${J_{i}}=-\overline{D}{\nabla_{i}}\rho+{\nabla_{j}}\sigma_{ij}$, where $\sigma_{ij}\equiv\alpha\nabla_{i}h\nabla_{j}h+\beta e_{jm}\nabla_{i}h\nabla_{m}h$ is reminiscent of “active stresses” found in active matter theories Marchetti et al. (2013), the $\beta$-term is a chiral contribution. The quadratic dependence of $\bf J$ on ${\bm{\nabla}}h$ implies the active particles (i) respond, unsurprisingly, to the height fluctuations, but not the absolute height, and (ii) ignoring gravity, the particles do not distinguish valleys from the hills (although the surface itself breaks the inversion symmetry). Here, $\overline{D}>0$ is a diffusivity. We focus on the quasi-static limit of infinitely fast dynamics of $\rho$, such that $\partial\rho/\partial t\approx 0$, giving $\overline{D}\nabla^{2}\rho=\alpha\nabla_{i}\nabla_{j}(\nabla_{i}h\nabla_{j}h)+\beta e_{jm}\nabla_{i}\nabla_{j}(\nabla_{i}h\nabla_{m}h)$ neglecting any noise in the $\rho$-dynamics. Now use this to eliminate $\rho$ in (2) to get (15), after absorbing a factor of $\overline{D}$. (We have implicitly assumed $\alpha,\beta$ to scale with $\overline{D}$, and ignored any advective-type nonlinearity originating from projecting the particle dynamics on the plane of the membrane in the large $\overline{D}$-limit.) All of $\lambda,\lambda_{1},\lambda_{2}$ can be individually positive and negative. The chiral term is 2D specific; the other two nonlinear terms with coefficients $\lambda$ and $\lambda_{1}$ can exist in any dimension $d$. Thus the $\lambda_{1}$- and $\lambda_{2}$-terms in (15) are physical, although our active species origin need not be the only possible source of these two terms. See Ref. Rasshofer et al. (2022) for a similar mechanism to generate an effective nonlocal dynamics in the noisy Fisher-Kolmogorov equation FISHER (1937); Kolmogorov et al. (1937) for population dynamics coupled with a fast chemical signal. Further Eq. (15) can be realised microscopically by considering an “active” 2D single-step model for a 2D KPZ surface with point particles living on it. The dynamical update rules of the modified single-step model now depend on the local excess or deficit population of the active particles, instead of being constants as they are in standard single-step models Meakin et al. (1986); Plischke et al. (1987); Liu and Plischke (1988); Kondev et al. (2000); Naserabadi et al. (2017). The particle hopping rates to the nearest neighbour sites in turn depends not only on the number inhomogeneities, but also on the height fluctuations (but without distinguishing local valleys from hills). Monte-Carlo simulations of this model, focusing on the limiting case of fast dynamics by the number fluctuations, should bring out the physics described in this Letter. The limit of fast particle dynamics can be implemented by considering time-scale separations in the rates of particle position updates and surface conformation updates.
@@
At one dimension, the $\lambda_{2}$-term vanishes, and the $\lambda_{1}$-term becomes indistinguishable from the $\lambda$-term. The transformation $x_{i}^{\prime}=x_{i}-(\lambda+2\lambda_{1})c_{i}t-\lambda_{2}e_{ij}c_{j}t$ and $t^{\prime}=t$, together with the height function $h$ transforming as $h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+\textbf{c}\cdot\textbf{x}$ leaves Eq. (15) invariant; see SM. This generalises invariance of the usual KPZ equation under a pseudo-Galilean transformation Barabási and Stanley (1995).
Similar to the KPZ equation, dimensional analysis via scaling ${\bf r}\rightarrow b{\bf r},\,t\rightarrow b^{z}t,\,h\rightarrow b^{\chi}h$, where $z$ and $\chi$ are the dynamic and roughness exponents, all of $\lambda,\lambda_{1},\lambda_{2}$ scale similarly, and $d=2$ is the critical dimension of Eq. (15); see SM. Whether it is the upper or lower critical dimension requires further analysis that follows below.
@@
We first determine if Eq. (15) has a stable nonequilibrium steady state (NESS), and second, if so, the scaling properties in those NESS. We use renormalisation group (RG) framework, well-suited to systematically handle the diverging corrections encountered in naïve perturbation theories. The dynamic RG method for our model closely resembles that for the KPZ equation Kardar et al. (1986); Barabási and Stanley (1995); Täuber (2014).
We perform the Wilson momentum shell dynamic RG procedure at the one-loop order Barabási and Stanley (1995); Täuber (2014); Forster et al. (1977); see SM for the one-loop Feynman diagrams. There are no one-loop corrections to $\lambda,\lambda_{1},\lambda_{2}$. However, there are diverging one-loop corrections to $\nu$ and $D$. Dimensional analysis allows us to identify an effective dimensionless coupling constant $g$ and two dimensionless ratios $\gamma_{1},\,\gamma_{2}$ defined as $g\equiv\frac{\lambda^{2}D}{\nu^{3}}\frac{1}{2\pi},\,\gamma_{1}\equiv\frac{\lambda_{1}}{\lambda},\,\gamma_{2}\equiv\frac{\lambda_{2}}{\lambda}$. By following the standard steps of RG, we obtain the RG recursion relations for $D,\,\nu$ at the one-loop order (here $l$ is the “RG time”; $\exp(l)$ is a length-scale):
@@
with $\gamma_{1},\,\gamma_{2}$ being marginal at the one-loop order, stemming from the nonrenormalisation of $\lambda,\lambda_{1},\lambda_{2}$ at that order. Equations (3)-(4) are invariant under the inversion of $\lambda_{2}$, i.e., interchange of left handed coordinate system with a right handed one.Here, ${\cal B}[\gamma_{1},\gamma_{2}]=\frac{3}{8}\gamma_{1}^{2}+\frac{1}{2}\gamma_{1}+\frac{1}{8}\gamma_{2}^{2}+\frac{1}{4},\,{\cal C}[\gamma_{1},\gamma_{2}]=\frac{1}{2}\gamma_{1}^{2}+\frac{5}{8}\gamma_{1}+\frac{1}{8}\gamma_{2}^{2}$; ${\cal B}>0$. Flow equations (3)-(4) yield the flow equation for $g$:
@@
We first focus on the achiral case, i.e., $\gamma_{2}=0$. An RG flow diagram in the $g-\gamma_{1}$ plane is shown in Fig. 1(a). The condition ${\cal\tilde{A}}(\gamma_{1})\equiv{\cal A}(\gamma_{1},\gamma_{2}=0)=0$ defines two solid (black) lines $\gamma_{1}=\gamma_{+},\,\gamma_{-}$ parallel to the $g$-axis in the $g-\gamma_{1}$ plane, where $\gamma_{+}=0.161,\gamma_{-}=-1.383$, such that for $\gamma_{+}>\gamma_{1}>\gamma_{-}$ (gray region), the RG flow lines flow away parallel to the $g$-axis towards infinity, indicating a perturbatively inaccessible, presumably rough, phase with short-ranged positional order. In this unstable region $g(l)$ diverges as $l\rightarrow 1/[|{\cal\tilde{A}}(\gamma_{1})|]$, reminiscent of the 2D KPZ equation Barabási and Stanley (1995), presumably corresponding to algebraically rough phase Basu and Frey (2004); Gomes-Filho et al. (2021). Outside this region, where ${\cal\tilde{A}}(\gamma_{1})>0$, the flow lines flow towards $g=0$ parallel to the $g$-axis, implying stability, with $g(l)\approx 1/[l{\cal\tilde{A}}(\gamma_{1})]$ vanishes slowly in the long wavelength limit $l\rightarrow\infty$. Although $g^{*}=0$ is the only fixed point (FP) in the stable region, the vanishing of $g(l)$ is so slow, being proportional to $1/l$, the parameters $D$ and $\nu$ are infinitely renormalised, altering the linear theory scaling in the long wavelength limit. The simplest way to see this is to set $z=2,\,\chi=0$ (i.e., their linear theory values) in (3) and (4) with $\gamma_{2}=0$, which gives
@@
Figure 1: (a) RG flow diagram in the $g-\gamma_{1}$ plane in the achiral limit ($\gamma_{2}=0$). Arrows indicate RG flows. Flow in the stable (unstable), i.e., towards (away from) $g=0$ region are marked. (b) Variations of $\mu$ and $\kappa$ as functions of $\gamma_{1}$ in the stable region for the achiral case. (c) RG flow diagram in the space spanned by $\gamma_{1}-\gamma_{2}-g$ in the full a-KPZ equation. RG flow lines in the stable and unstable regions are shown by the arrows. (d) Phase diagram in the $\gamma_{1}-\gamma_{2}$ plane for the a-KPZ equation. The central gray region containing the origin is unstable. Regions with SQLRO and WQLRO are marked (see text).
@@
The chiral effects ($\gamma_{2}\neq 0$) can be included in the above calculational scheme in straightforward ways: Stability of the RG flow is now determined by ${\cal A}(\gamma_{1},\gamma_{2})>0$. Flow lines having initial conditions within a narrow elliptical cylinder, containing the origin (0,0,0) and having the axis parallel to the $g$-axis, with its surface given by ${\cal A}(\gamma_{1},\gamma_{2})=0$ for any $g$, run away parallel to the $g$-axis, leaving the perturbatively accessible region. Flow lines with initial conditions falling in regions outside of this elliptical cylinder flow towards the $\gamma_{1}-\gamma_{2}$ plane with stable states. See Fig. 1(c) depicting the RG flow lines in the space spanned by $\gamma_{1}-\gamma_{2}-g$. Outside the elliptical cylinder $g(l)\sim 1/({\cal A}l)$ for large $l$, similar to its achiral analog. Inside the cylinder, $g(l)$ diverges as $l\rightarrow 1/(|{\cal A}|)$ from below. Focusing on the $\gamma_{1}-\gamma_{2}$ plane, ${\cal A}(\gamma_{1},\gamma_{2})=0$ sketches out an inner elliptical unstable region, whereas the outer region is stable; see Fig.1(d). We use the above results to find that in the stable region $\Delta\sim[\ln(L/a)]^{\mu(\gamma_{1},\gamma_{2})}$, where $\mu=1+({\cal B}-{\cal C})/{\cal A}$ is now parametrised by both $\gamma_{1},\gamma_{2}$. Similar to and quantitatively extending the achiral case, $\mu<1(>1)$ is referred to as SQLRO (WQLRO), giving positional generalised QLRO. The SQLRO and WQLRO regions are demarcated within the stable region in Fig. 1(d).

### v4

@@ Abstract
The Kardar-Parisi-Zhang (KPZ) equation sets the universality class for growing and roughening of nonequilibrium surfaces without any conservation law and nonlocal effects. We argue here that the KPZ equation can be generalized by including a symmetry-permitted nonlocal nonlinear term of active origin that is of the same order as the one included in the KPZ equation. Including this term, the 2D active KPZ equation is stable in some parameter regimes, in which the interface conformation fluctuations exhibit sublogarithmic or superlogarithmic roughness, with nonuniversal exponents, giving positional generalised quasi-long-ranged order. For other parameter choices, the model is unstable, suggesting a perturbatively inaccessible algebraically rough interface or positional short-ranged order. Our model should serve as a paradigmatic nonlocal growth equation.
@@
The Kardar-Parisi-Zhang (KPZ) equation Kardar et al. (1986); Krug (1997); Barabási and Stanley (1995) for growing nonequilibrium surfaces displays a nonequilibrium roughening transition between a smooth phase, whose long wavelength scaling properties are identical to an Edward-Wilkinson (EW) surface Edwards and Wilkinson (1982), to a perturbatively inaccessible rough surface Barabási and Stanley (1995); Tang et al. (1990) when $d>d_{c}=2$, its lower critical dimension. Importantly, the local normal velocity of a KPZ surface depends locally on surface fluctuations, and hence cannot describe nonequilibrium surface dynamics with nonlocal interactions.
@@
Theoretical studies on nonlocal interactions has a longstanding history in equilibrium systems Fisher et al. (1972); Frey and Schwabl (1994); Bayong et al. (1999); Aronovitz and Lubensky (1988); Banerjee et al. (2019a, b). Examples of their prominent nonequilibrium counterparts include interface dynamics involving nonlocal interactions, e.g., flame front propagation, thin film growth Pelcé et al. (2004), and shading phenomena in surface growth Ketterl and Hagston (1999). Kinetic roughening in the presence of nonlocal interactions Mukherji and Bhattacharjee (1997) display generic non-KPZ scaling behavior. Nonlocal effects are often important in biological growth processes; see, e.g., Ref. Santalla and Ferreira (2018) for a recent study. Furthermore, in many applications, the growth is controlled by fast nonlocal transport not included in the KPZ equation. Prominent examples include diffusion-controlled nonlocal transport Aharonov and Rothman (1996), dissolution or precipitation processes Lichtner et al. (1996), gas-solid reactions Doǧu (1981), a variety of reaction engineering processes Delmono and Froment (1999), diffusion-limited erosion, that displays nonlocal stabilization of surfaces Krug and Meakin (1991) (see also Ref. Nicoli et al. (2009)), and even geological contexts, e.g., earth surface roughness R. Schumer (2017). Inspired by these past studies, we explore the generic consequences of competition between local contributions and those that depend on the global surface profile, i.e., nonlocal contributions to the local surface velocity, by constructing a purpose-built conceptual model.
@@
In this Letter, we set up and study a generalization of the KPZ equation, where the surface velocity depends, in contrast to the KPZ equation, nonlocally on the surface fluctuations. We do this by adding symmetry-permitted nonlocal nonlinear gradient terms that are of the same order as the usual KPZ nonlinear term. These nonlocal, nonlinear terms have the same scaling as the usual local nonlinear term of the KPZ equation. This allows us to study competition and interplay between local and nonlocal nonlinear effects, resulting into stable steady states and roughening transitions distinct from both the usual KPZ equation, or the KPZ equation with truly long-range effects (with either long range nonlinearity or long range noises) Medina et al. (1989); Mukherji and Bhattacharjee (1997). To generalize the scope of our study, we also include chiral contributions, which is ubiquitous in soft matter and biologically inspired systems; see, e.g., Refs. Martin et al. (2021); Lalitha et al. (2001). The resulting equation in 2D, named active-KPZ or a-KPZ equation, is
@@
a nonlocal generalization to the usual KPZ equation that is distinct from the one considered in Ref. Mukherji and Bhattacharjee (1997). Here, the tensor $e_{jm}$ is the 2D totally antisymmetric matrix. Further, $Q_{ij}({\bf r})$ is the longitudinal projection operator that in the Fourier space is $Q_{ij}({\bf k})=k_{i}k_{j}/k^{2}$, where $\bf k$} is a Fourier wave vector, and is nonlocal. Physically, $\lambda_{1}Q_{ij}({\bf r})(\nabla_{i}h\nabla_{j}h)+\lambda_{2}Q_{ij}({\bf r})e_{jm}(\nabla_{i}h\nabla_{m}h)$ is the contribution to the surface velocity normal to the base plane $v_{p}=\partial h/\partial t$ that is nonlocal in height fluctuations ${\bm{\nabla}}h$. Noise $\eta$ is a zero-mean, Gaussian-distributed white noise with a variance $\langle\eta({\bf x},t)\eta({\bf 0},0)\rangle=2D\delta^{d}({\bf x})\delta(t)$. We extract the scaling of the stable phases, which exists for a range of the model parameters. In particular, we show that the variance $\Delta\equiv\langle h({\bf x},t)^{2}\rangle\sim[\ln(L/a)]^{\mu}$ for a surface of lateral size $L$, where $\mu<(>)1$ for sub (super) logarithmic roughness and $a$ is a microscopic cutoff. This defines positional generalized quasi-long-ranged order (QLRO), generalizing the well-known QLRO of EW surfaces Barabási and Stanley (1995), in which $\Delta\sim\ln(L/a)$, i.e., $\mu=1$. Further, the time-scale of relaxation $\tau(L)\sim L^{2}\,(\ln\,L)^{-\kappa},\,\kappa>0$, i.e., logarithmically superdiffusive. Both $\mu$ and $\kappa$ are nonuniversal. They vary continuously with $\lambda_{1}/\lambda,\,\lambda_{2}/\lambda$.
The form of Eq. (1) can be obtained by first considering the mapping from the KPZ equation to the Burgers equation Forster et al. (1977) in terms of the “Burgers velocity ${\bf v}={\bm{\nabla}}h$.” Now generalizing the Burgers equation nonlinearity $\lambda{\bm{\nabla}}v^{2}$ to $\lambda_{1}\nabla_{j}(v_{i}v_{j})+\lambda_{2}e_{jm}\nabla_{j}(v_{i}v_{m})$, and then writing them in terms of $h$ produces the $\lambda_{1}$ and $\lambda_{2}$ terms in (1); see the Supplemental Material (SM) sup .
The $\lambda_{1}$ and $\lambda_{2}$ terms in (1) can be motivated by considering a nearly flat nonequilibrium surface without any momentum conservation described by a single valued height field $h({\bf x},t)$ in Monge gauge  Chaikin and Lubensky (2000); Nelson et al. (1989), with an active conserved density $\rho({\bf x},t)$ living on it. Its hydrodynamic equation, retaining only the lowest order in nonlinearities and spatial gradients, reads
@@
where $v(\rho)$ is a local density-dependent velocity of the membrane; $v(\rho)=v_{0}+g_{1}\rho$ to the leading order in $\rho$; $g_{1}$ is a coupling constant of either sign. Further, density $\rho$ follows $\partial_{t}\rho=-{\bm{\nabla}}\cdot{\bf J}$, where $\bf J$ is the current. The specific form of the particle dynamics decides the structure of $\bf J$. We choose ${J_{i}}=-\overline{D}{\nabla_{i}}\rho+{\nabla_{j}}\sigma_{ij}$, where $\sigma_{ij}\equiv\alpha\nabla_{i}h\nabla_{j}h+\beta e_{jm}\nabla_{i}h\nabla_{m}h$ is reminiscent of “active stresses” found in active matter theories Marchetti et al. (2013), the $\beta$ term is a chiral contribution. The quadratic dependence of $\bf J$ on ${\bm{\nabla}}h$ implies the active particles (i) respond, unsurprisingly, to the height fluctuations, but not the absolute height; and (ii) ignoring gravity, the particles do not distinguish valleys from the hills (although the surface itself breaks the inversion symmetry). Here, $\overline{D}>0$ is a diffusivity. We focus on the quasistatic limit of infinitely fast dynamics of $\rho$, such that $\partial\rho/\partial t\approx 0$, giving $\overline{D}\nabla^{2}\rho=\alpha\nabla_{i}\nabla_{j}(\nabla_{i}h\nabla_{j}h)+\beta e_{jm}\nabla_{i}\nabla_{j}(\nabla_{i}h\nabla_{m}h)$ neglecting any noise in the $\rho$ dynamics. Now use this to eliminate $\rho$ in (2) to get (1), after absorbing a factor of $\overline{D}$. (We have implicitly assumed $\alpha,\beta$ to scale with $\overline{D}$, and ignored any advective-type nonlinearity originating from projecting the particle dynamics on the plane of the membrane in the large $\overline{D}$ limit). All of $\lambda,\lambda_{1},\lambda_{2}$ can be individually positive and negative. The chiral term is 2D specific; the other two nonlinear terms with coefficients $\lambda$ and $\lambda_{1}$ can exist in any dimension $d$. Thus the $\lambda_{1}$ and $\lambda_{2}$ terms in (1) are physical, although our active species origin need not be the only possible source of these two terms. See Ref. Kolk et al. (2023) for a similar mechanism to generate an effective nonlocal dynamics in the noisy Fisher-Kolmogorov equation FISHER (1937); Kolmogorov et al. (1937) for population dynamics coupled with a fast chemical signal. Further Eq. (1) can be realized microscopically by considering an “active” 2D single-step model for a 2D KPZ surface with point particles living on it. The dynamical update rules of the modified single-step model now depend on the local excess or deficit population of the active particles, instead of being constants as they are in standard single-step models Meakin et al. (1986); Plischke et al. (1987); Liu and Plischke (1988); Kondev et al. (2000); Naserabadi et al. (2017). The particle hopping rates to the nearest neighbor sites in turn depends not only on the number inhomogeneities, but also on the height fluctuations (but without distinguishing local valleys from hills). Monte-Carlo simulations of this model, focusing on the limiting case of fast dynamics by the number fluctuations, should bring out the physics described in this Letter. The limit of fast particle dynamics can be implemented by considering time-scale separations in the rates of particle position updates and surface conformation updates.
@@
At one dimension, the $\lambda_{2}$ term vanishes, and the $\lambda_{1}$ term becomes indistinguishable from the $\lambda$-term. The transformation $x_{i}^{\prime}=x_{i}-(\lambda+2\lambda_{1})c_{i}t-\lambda_{2}e_{ij}c_{j}t$ and $t^{\prime}=t$, together with the height function $h$ transforming as $h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+\textbf{c}\cdot\textbf{x},$ leaves Eq. (1) invariant; see the SM sup . This generalizes invariance of the usual KPZ equation under a pseudo-Galilean transformation Barabási and Stanley (1995).
Similar to the KPZ equation, dimensional analysis via scaling ${\bf r}\rightarrow b{\bf r},\,t\rightarrow b^{z}t,\,h\rightarrow b^{\chi}h$, where $z$ and $\chi$ are the dynamic and roughness exponents, reveals that all of $\lambda,\lambda_{1},\lambda_{2}$ scale similarly, and hence are equally relevant (in the scaling sense). Furthermore, $d=2$ is the critical dimension of Eq. (1); see the SM sup . Whether it is the upper or lower critical dimension requires further analysis that follows below. That all of $\lambda,\lambda_{1},\lambda_{2}$ scale the same way is important: it means the nonlocal, nonlinear effects in (1) are as relevant as the short-range, local nonlinear effects in the original KPZ equation Barabási and Stanley (1995). This feature clearly distinguishes the active KPZ equation (1) from generalized KPZ equations with genuine long-range interactions Mukherji and Bhattacharjee (1997). Indeed, just as the usual KPZ equation Barabási and Stanley (1995) is universal in the sense that all short-range growth processes with just one soft mode (height $h$) and without any conservation laws, inversion symmetry and disorder should be described by it; the active KPZ equation (1) should likewise describe all such nonlocal growth processes having the same scaling properties as the corresponding local growth processes in the KPZ equation, and with just one soft mode ($h$) but without any conservation laws, inversion symmetry, and disorder, highlighting the universal nature of (1).
@@
We first determine if Eq. (1) has a stable nonequilibrium steady state (NESS), and second, if so, the scaling properties in those NESS. We use renormalization group (RG) framework, well suited to systematically handle the diverging corrections encountered in naïve perturbation theories. The Wilson dynamic RG method for our model closely resembles that for the KPZ equation Kardar et al. (1986); Barabási and Stanley (1995); Forster et al. (1977); Täuber (2014); see the SM sup for the one-loop Feynman diagrams. There are no one-loop corrections to $\lambda,\lambda_{1},\lambda_{2}$. However, there are diverging one-loop corrections to $\nu$ and $D$. Dimensional analysis allows us to identify an effective dimensionless coupling constant $g$ and two dimensionless ratios $\gamma_{1},\,\gamma_{2}$ defined as $g\equiv\frac{\lambda^{2}D}{\nu^{3}}\frac{1}{2\pi},\,\gamma_{1}\equiv\frac{\lambda_{1}}{\lambda},\,\gamma_{2}\equiv\frac{\lambda_{2}}{\lambda}$. The RG recursion relations for $D,\,\nu$ at the one-loop order (here $l$ is the “RG time;” $\exp(l)$ is a length scale)
@@
with $\gamma_{1},\,\gamma_{2}$ being marginal at the one-loop order, stemming from the nonrenormalization of $\lambda,\lambda_{1},\lambda_{2}$ at that order. Here, ${\cal B}[\gamma_{1},\gamma_{2}]=\frac{3}{8}\gamma_{1}^{2}+\frac{1}{2}\gamma_{1}+\frac{1}{8}\gamma_{2}^{2}+\frac{1}{4},\,{\cal C}[\gamma_{1},\gamma_{2}]=\frac{1}{2}\gamma_{1}^{2}+\frac{5}{8}\gamma_{1}+\frac{1}{8}\gamma_{2}^{2}$; ${\cal B}>0$. Flow equations (3) and (4) yield the flow equation for $g$:
@@
In the achiral case, i.e., $\gamma_{2}=0$, an RG flow diagram in the $g$-$\gamma_{1}$ plane is shown in Fig. 1(a). The condition ${\cal\tilde{A}}(\gamma_{1})\equiv{\cal A}(\gamma_{1},\gamma_{2}=0)=0$ defines two solid (black) lines $\gamma_{1}=\gamma_{+},\,\gamma_{-}$ parallel to the $g$ axis in the $g$-$\gamma_{1}$ plane, where $\gamma_{+}=0.161,\gamma_{-}=-1.383$, such that for $\gamma_{+}>\gamma_{1}>\gamma_{-}$ (gray region), the RG flow lines flow away parallel to the $g$ axis toward infinity, indicating a perturbatively inaccessible, presumably rough, phase with short-ranged positional order. In this unstable region $g(l)$ diverges as $l\rightarrow 1/[|{\cal\tilde{A}}(\gamma_{1})|]$, reminiscent of the 2D KPZ equation Barabási and Stanley (1995), presumably corresponding to algebraically rough phase Basu and Frey (2004); Gomes-Filho et al. (2021). Outside this region, where ${\cal\tilde{A}}(\gamma_{1})>0$, the flow lines flow toward $g=0$ parallel to the $g$ axis, implying stability, while $g(l)\approx 1/[l{\cal\tilde{A}}(\gamma_{1})]$ vanishes slowly in the long wavelength limit $l\rightarrow\infty$. Although $g^{*}=0$ is the only fixed point (FP) in the stable region, the vanishing of $g(l)$ is so slow, being proportional to $1/l$, the parameters $D$ and $\nu$ are infinitely renormalized, altering the linear theory scaling in the long wavelength limit. The simplest way to see this is to set $z=2,\,\chi=0$ (i.e., their linear theory values) in (3) and (4) with $\gamma_{2}=0$, which gives
@@
Figure 1: (a) RG flow diagram in the $g$-$\gamma_{1}$ plane in the achiral limit ($\gamma_{2}=0$). Arrows indicate RG flows. Flow in the stable (unstable), i.e., toward (away from), $g=0$ region are marked. (b) Variations of $\mu$ and $\kappa$ as functions of $\gamma_{1}$ in the stable region for the achiral case. (c) RG flow diagram in the space spanned by $\gamma_{1}$-$\gamma_{2}$-$g$ in the full a-KPZ equation. RG flow lines in the stable and unstable regions are shown by the arrows. (d) Phase diagram in the $\gamma_{1}$-$\gamma_{2}$ plane for the a-KPZ equation. The central gray region containing the origin is unstable. Regions with SQLRO and WQLRO are marked (see text).
@@
Including the chiral effects ($\gamma_{2}\neq 0$), stability of the RG flow is now determined by ${\cal A}(\gamma_{1},\gamma_{2})>0$. Flow lines having initial conditions within a narrow elliptical cylinder, containing the origin (0,0,0), and having the axis parallel to the $g$ axis, with its surface given by ${\cal A}(\gamma_{1},\gamma_{2})=0$ for any $g$, run away parallel to the $g$ axis, leaving the perturbatively accessible region. Flow lines with initial conditions falling in regions outside of this elliptical cylinder flow toward the $\gamma_{1}$-$\gamma_{2}$ plane with stable states. See Fig. 1(c) depicting the RG flow lines in the space spanned by $\gamma_{1}$-$\gamma_{2}$-$g$. Outside the elliptical cylinder $g(l)\sim 1/({\cal A}l)$ for large $l$, similar to its achiral analog. Inside the cylinder, $g(l)$ diverges as $l\rightarrow 1/(|{\cal A}|)$ from below. Focusing on the $\gamma_{1}$-$\gamma_{2}$ plane, ${\cal A}(\gamma_{1},\gamma_{2})=0$ sketches out an inner elliptical unstable region, whereas the outer region is stable; see Fig.1(d). We use the above results to find that in the stable region $\Delta\sim[\ln(L/a)]^{\mu(\gamma_{1},\gamma_{2})}$, where $\mu=1+({\cal B}-{\cal C})/{\cal A}$ is now parametrized by both $\gamma_{1},\gamma_{2}$. Similar to and quantitatively extending the achiral case, $\mu<1(>1)$ is referred to as SQLRO (WQLRO), giving positional generalised QLRO. The SQLRO and WQLRO regions are demarcated within the stable region in Fig. 1(d).